OpenStudy (anonymous):

Help I am trying to study for a test.... given the following function f(x)=2x^2-4x-5\ 1) determine the vertex 2) list the axis of symmetry 3) calculate the y-intercept 4) use the axis of symmetry and the y-intercept to find an additional point on the graph

OpenStudy (mertsj):

Do you know that it's a parabola?

OpenStudy (anonymous):

No that is all the study paper give us was this ? exactly like it is

OpenStudy (mertsj):

What have you been studying in your class?

OpenStudy (anonymous):

This is the start of the class and we are having to take a pretest

OpenStudy (anonymous):

it is college algebra or precalculus

OpenStudy (mertsj):

well then your teacher is trying to find out what you know so just do the best you can.

OpenStudy (anonymous):

well out textbook does not really give a good explanation of anything

OpenStudy (mertsj):

So do you want to work through this problem?

OpenStudy (anonymous):

i would appreciate any help i can get

OpenStudy (mertsj):

This is the equation of a parabola. The x coordinate of the vertex can be found using the formula x = -b/2a

OpenStudy (mertsj):

which in this case is 4/4 or 1

OpenStudy (mertsj):

Now you can find the y coordinate of the vertex by replacing x with 1 and finding f(x)

OpenStudy (mertsj):

f(1)=-7 So the vertex is (1,-7)

OpenStudy (anonymous):

so how do u come up with the 7

OpenStudy (mertsj):

The axis of symmetry is the vertical line through the vertex so it is x = 1

OpenStudy (mertsj):

Did you replace x with 1 in your equation?

OpenStudy (anonymous):

o no

OpenStudy (mertsj):

\[f(x)=2(1)^2-4(1)-5=2-4-5=2-9=-7\]

OpenStudy (mertsj):

Are you reading what I write?

OpenStudy (mertsj):

Now you can find the y coordinate of the vertex by replacing x with 1 and finding f(x)

OpenStudy (anonymous):

yes I am

OpenStudy (mertsj):

So now we are ready for 3 which is calculate the y intercept. We can find the y intercept by letting x = 0. So replace x with 0

OpenStudy (mertsj):

Did you get -5? So the y intercept is (0,-5)

OpenStudy (anonymous):

just a sec

OpenStudy (mertsj):

one second granted

OpenStudy (anonymous):

ok 0 cancels so i end up with -5

OpenStudy (mertsj):

Yep

OpenStudy (mertsj):

Now we're ready for the last part. And for that we need to put the info we have on a graph.

OpenStudy (mertsj):